The television market in South Africa has two major suppliers producing differentiated television models. The monthly demand functions of the two suppliers is as follows:

Samsung q1= 450 - 2p1 + p2
LG q2 = 450 - 2p2 + p1

The marginal costs of production are R60 per unit and fixed costs are equal to zero

Part 1
a) write down an expression for Samsungs inverse demand function which expresses p1 on terms of q1 and q2 only

b) in a similar manner, write down the expression of LGs inverse demand function in the form of an equation for p2 in terms of q1 and q2 only
1
Expert's answer
2020-07-17T12:57:02-0400

(a)

"p1=q2\u2212450+2(q1\u2212450+2p1)"

"p1=q2\u2212450+2q1\u2212900+4p1"

"3p1=1350\u2212q2\u22122q1"

"p1 = 450 - \frac{1}{3} q2 - \frac{2}{3}q1"


(b)

"p2 = q1-450+2(q2-450-2p2)"

"p2=q1\u2212450+2q2\u2212900\u22124p2"

"3p2 = 1350 - q1 - q2"

"p2 = 450 - \frac{1}{3}q1 - \frac{2}{3} q2"